### v1

# ClaimGen-CN: A Large-scale Chinese Dataset for Legal Claim Generation

Abstract
Legal claims refer to the plaintiff’s demands in a case and are essential to guiding judicial reasoning and case resolution. While many works have focused on improving the efficiency of legal professionals, the research on helping non-professionals (e.g., plaintiffs) remains unexplored. This paper explores the problem of legal claim generation based on the given case’s facts. First, we construct ClaimGen-CN, the first dataset for Chinese legal claim generation task, from various real-world legal disputes. Additionally, we design an evaluation metric tailored for assessing the generated claims, which encompasses two essential dimensions: factuality and clarity. Building on this, we conduct a comprehensive zero-shot evaluation of state-of-the-art general and legal-domain large language models. Our findings highlight the limitations of the current models in factual precision and expressive clarity, pointing to the need for more targeted development in this domain. To encourage further exploration of this important task, we will make the dataset publicly available.
ClaimGen-CN: A Large-scale Chinese Dataset for Legal Claim Generation
Siying Zhou1, Yiquan Wu1, Hui Chen1, Xavier Hu1, Kun Kuang1∗ Adam Jatowt2, Ming Hu1, Chunyan Zheng1∗, Fei Wu1 1Zhejiang University, Hangzhou, China 2University of Innsbruck, Innsbruck, Austria {zhousiying, wuyiquan, 22402119, kunkuang,hm606, boxzheng}@zju.edu.cn xavier.hu.research@gmail.com, adam.jatowt@uibk.ac.at, wufei@cs.zju.edu.cn
11footnotetext: Corresponding Authors.
1 Introduction
Over the past decades, the advancement of natural language processing (NLP) techniques has advanced the field of Legal Artificial Intelligence (Legal AI). Legal AI is an important subfield of artificial intelligence, with the goal of supporting individuals across various legal tasks, including legal judgment prediction (Zhong et al., 2018; Xu et al., 2020; Wu et al., 2023), court view generation (Wu et al., 2020), similar case matching (Bhattacharya et al., 2020), legal language understanding (Chalkidis et al., 2022), and legal question answering (Zhong et al., 2020b). As Legal AI systems become more advanced, it is essential to reflect not only on what these systems can do, but also on what they ought to contribute to society. If we could push the boundaries of what we envision AI to be, to the point where we explicitly require it to have a “positive impact on people and communities,” ensuring our definition of success explicitly includes that, AI could make the world a better place (Li, 2023). Yet, the existing research predominantly targets courtroom trials and judge assistance (Ma et al., 2021; Malik et al., 2021; Feng et al., 2022; Zhang et al., 2023; Le et al., 2024; Li et al., 2025), with limited attention to pre-trial contexts or non-professional needs, such as legal claim generation.
Figure 1: Conceptual overview of the differences between the pre-court scenario (left part) and the in-court scenario (right). In the pre-court scenario, claims are generated and prepared, which are then addressed and resolved during the in-court scenario. It is important to note that unreasonable claims may lead to a loss.
To promote the rule of law and make legal support more accessible, we advocate exploring the problem of legal claim generation. As Figure 1 shows, different legal scenarios play different roles. The pre-court scenario is primarily dedicated to preparing claims for the plaintiffs, whereas the in-court scenario is where those claims are utilized. Our work is concentrated on the pre-court phase, focusing on the interests of the parties involved. To the best of our knowledge, we are the first to explore the problem of legal claim generation and to reaffirm its significance in civil litigation.
Technically, legal claim generation introduces two key challenges that set it apart from prior Legal AI tasks. First, it is inherently open-ended: models must generate legal claims directly from factual narratives, without the guidance of predefined templates, as is common in judgment generation. Second, the input is typically authored by non-experts. Unlike court view generation, which builds on legally structured texts written by judges, this task requires interpreting informal, often unstructured, and emotional descriptions. The model must extract relevant facts, infer the underlying legal intent, and articulate it as a clear and valid claim.
To support research on legal AI in the field of civil litigation, several legal datasets have been constructed (Wang et al., 2018; Long et al., 2019; Xiao et al., 2021). However, existing work still faces the following challenges: 1) Insufficient Data Coverage: Many existing datasets are limited to one or a few causes of action. For instance, both AC-NLG (Wu et al., 2020) and MSJudge (Ma et al., 2021) focus only on the private lending category in civil cases. 2) Lack of diverse and fine-grained metrics: Commonly used metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) focus mainly on n-gram overlap and are not comprehensive enough for evaluating the completeness of key information in legal texts.
To address the limited scope and diversity of existing legal datasets, we manually construct a dataset from real-world cases covering one hundred causes of action. Moreover, to better evaluate model performance in claim generation, we introduce several fine-grained criteria, which include two essential dimensions: factuality and clarity. To better understand current model capabilities, we further conduct a comprehensive zero-shot evaluation of state-of-the-art general and legal-domain large language models on the legal claim generation task.
To summarize, our main contributions are as follows:
  • We explore legal AI from the perspective of non-professionals (e.g., plaintiffs), and extend its scope to a novel and significant problem: legal claim generation.
  • We construct a large-scale Chinese dataset, ClaimGen-CN, built from raw civil legal documents. Furthermore, we introduce two key aspects for assessing claims: factuality and clarity.
  • We conduct a comprehensive zero-shot evaluation of state-of-the-art general and legal-domain large language models on the legal claim generation task. Our detailed error analysis reveals notable limitations in current models, particularly in terms of factual accuracy and linguistic clarity.
  • To motivate other scholars to investigate this novel and important problem, we make the ClaimGen-CN dataset publicly available [^1] https://github.com/JosieZhou00/ClaimGen-CN.
2 Related Work
2.1 Legal AI
Legal AI aims to enhance tasks within the legal domain through the utilization of artificial intelligence techniques (Zhong et al., 2020a; Katz et al., 2023). The field has evolved from foundational developments in legal question answering and entity recognition, as explored by early researchers like Monroy et al. (2009) and Cardellino et al. (2017), to more advanced applications. The integration of AI in areas such as court view generation and legal summarization has shown its potential to process and interpret complex legal texts (Wu et al., 2020; Hachey and Grover, 2006). Recent advancements in legal language understanding and predictive models for legal judgments demonstrate the significant strides made in applying AI to legal analysis (Bommarito et al., 2018; Zhong et al., 2018; Liu et al., 2023a). The ongoing innovations in the field, particularly in the development of legal pre-trained models and event detection in legal documents, are further evidenced in the research by Chalkidis et al. (2020), showcasing the transformative impact of AI on legal research and practice.
2.2 Legal Datasets
Recent advancements in Legal AI have been significantly driven by the development and use of specialized legal datasets. For instance, the European Court of Human Rights (ECHR) dataset provided rich text data for human rights case law research, further diversifying the applications of AI in legal research (Aletras et al., 2016). LEXGLUE dataset provided a comprehensive benchmark for various legal NLP tasks, offering an invaluable resource for testing and improving AI models in the legal domain (Chalkidis et al., 2020). Additionally, the Contract Understanding Atticus Dataset (CUAD) was designed for AI-based contract review and analysis, enabling more accurate and efficient processing of legal contracts (Hendrycks et al., 2021). TWLJP dataset was collected using indictments to assist prosecutors in indexing charges and crime factual descriptions (Chien et al., 2024).
However, existing works primarily aim to assist experts such as judges and prosecutors, often overlooking common individuals in need of help with regards to their infringed rights. Our work shifts the research perspective from aiding legal professionals to assisting the non-expert public.
3 Problem Formulation
We first clarify the definition of the terms as follows.
$\bullet$ Facts and Reasons consist of two parts: a brief narrative detailing the dispute’s background, including the legal relationship, time, place, specific case details, cause, process, circumstances, and consequences; and the grounds upon which the plaintiff seeks court relief. Evidence is encompassed in the plaintiff’s presentation of facts and reasons. We define facts and reasons as a token sequence $f$.
$\bullet$ Claims can be singular or multiple, each representing a specific demand to the court for the protection of civil rights. A claim is defined as a token sequence $c$. For a given case, there can be several claims, which can be denoted as $C=\{c_{1},c_{2},...,c_{n}\}$, where $n$ is the number of claims.
Common types of legal claims include: 1) Confirmation of Legal Relationships. This claim seeks court confirmation for certain legal relationships, such as the confirmation of the status of a missing or deceased individual. 2) Enforcement of Obligations. This claim compels the opposing party to fulfill payment obligations. Examples include seeking compensation for losses, demanding the repayment of loan principal and interest, or demanding compliance with contractual obligations. 3) Alteration or Termination of Civil Legal Relationships. This claim covers demands for changes or termination of specific civil legal relationships. Examples include filing for divorce or demanding cancellations of contracts.
Then the problem can be defined as:
Problem 1 (Claim generation).
Given the facts and reasons $f$, the task is to predict the claims $C$.
4 Dataset
4.1 Dataset Construction
We construct ClaimGen-CN, a Chinese dataset for legal claim generation, from 207,748 civil documents sourced from China Judgments Online [^2] http://wenshu.court.gov.cn/. While these documents span civil, criminal, and administrative cases, our work specifically focuses on civil cases for two main reasons.
$\bullet$ Civil cases, distinct from criminal or administrative cases, often involve private litigants and a wider variety of claims, making them more relevant for our research on claim generation. In criminal cases, since most claims are brought by professionals (prosecutors) and private prosecutions are rare (Krauss, 2009; Xiong, 2021), the generated claims carry a lesser significance in promoting the accessibility of judicial assistance. In administrative litigation, the court’s decisions are not strictly bound by the plaintiff’s claims (He, 2022); unlike civil litigation, which follows the “plaintiff requests, and the court adjudicates accordingly” approach, the generated claims do not have as significant an impact on judges in administrative cases as they do in civil cases.
$\bullet$ Civil cases constitute the majority of the three categories of cases, accounting for 87% of all cases, followed by criminal cases at 10%, and administrative cases at 3%, according to publicly available data [^3] The proportions of the three case types are calculated based on statistics from the China Judgments Online website as of December 6, 2023: 87,618,891 civil documents, 10,010,356 criminal documents, and 3,031,224 administrative documents.. Furthermore, civil litigation is the most common legal proceeding encountered in everyday life.
4.1.1 Raw Data Processing
Each initial document contains unstructured content over the entire document along with additional information, including the cause of action, title, start time, and end time. To maintain material homogeneity, we only select first-instance civil judgment documents and filter out those that are not publicly available for some reason. To ensure completeness, we segment the content, retaining only the task-related segments. To assess completeness, we use a sequential inclusion criterion with the following keywords: “file a lawsuit with this court”, “facts and reasons”, “argument”, “this court believes”, and “judgment as follows”. The content has been then segmented into specific sections, including introduction, plaintiff’s facts, plaintiff’s claims, defendant’s arguments, court’s findings, and judgment. For our task, we exclusively utilize plaintiff’s facts as input and plaintiff’s claims as output.
4.1.2 Structured Data
Among all the sampled data, there are 134 causes of action. Due to insufficient data in some categories, we retain the top 100 most common civil causes of action for our main dataset, which we call ClaimGen-CN. In addition, we constructed a test set, ClaimGen-CN-test, by exclusively selecting cases where the court fully supports the plaintiff’s claims. This allows us to obtain reference claims that are both reasonable and legally grounded. This test set still follows the original distribution of causes of action from the sampled data.
Table 1 presents the statistics of the processed datasets; all the experiments are conducted on the same datasets.
Type ClaimGen-CN ClaimGen-CN-test
# Cause of Action 100 100
# Sample 207,748 1,000
Avg. # Fact Tokens 353.4 379.2
Avg. # Claim Tokens 134.3 120.1
Table 1: Statistics of datasets.
Dataset # Samples # Causes Avg. # Claim Tokens Avg. # Fact Tokens Availability Plaintiff-centered
AutoJudge (Long et al., 2019) 100,000 1 23.9 100.1
AC-NLG (Wu et al., 2020) 66,904 1 77.9 158
MSJudge (Ma et al., 2021) 70,482 1 143
LK (Gan et al., 2021) 61,611 1
CCJudge (Zhao et al., 2021) 123,048 238 109.8 198.7
CPEE (Zhao et al., 2022) 158,625 10 99 156
ClaimGen-CN (Ours) 207,748 100 134.3 353.4
Table 2: Comparison of legal datasets. – means the value is not accessible.
4.2 Dataset Characteristics
The comparison of ClaimGen-CN with other open-access legal datasets is presented in Table 2. We discuss the characteristics of ClaimGen-CN based on the following aspects.
Diverse. Prior research mainly centered on private lending disputes, the most common cause of action. Among the existing open-source datasets, ClaimGen-CN stands out as the only one with more than 10 distinct civil case categories. ClaimGen-CN encompasses and openly shares a diverse range of case categories that were absent in prior work, including private lending disputes, divorce disputes, sales contract disputes, labor disputes, residential lease contract disputes, maintenance disputes, education and training contract disputes, and more.
Large-scale. As Table 2 illustrates, ClaimGen-CN is currently the largest civil litigation dataset, comprising 207k records and covering a wide variety of case types. This diversity in data provides researchers with numerous options for drafting legal claims. The number of tokens in the claims and facts in this dataset is much higher than in the previous datasets, which also introduces greater complexity.
Plaintiff-centered. ClaimGen-CN is a plaintiff-centered dataset that emphasizes the demands of plaintiffs. This dataset primarily focuses on the connection between the plaintiff’s facts and their claims. Additionally, it shifts the focus from supporting judges to helping the non-expert public whose rights have been violated.
5 Experiments
5.1 Baseline Models
We implement the following baseline models for comparison [^4] To our best knowledge, we are the first to propose the claim generation task, and there are no dedicated previous methods that could be directly applied for this task.: GPT-4o (Hurst et al., 2024) is an optimized version of GPT-4 (Achiam et al., 2023), featuring superior natural language processing capabilities and multimodal interaction functions. It can handle various data types including text, images, and audio, and is suitable for a wide range of complex tasks and application scenarios. LLaMA3.1 (Grattafiori et al., 2024) possesses powerful multilingual dialogue and code generation capabilities after pre-training and instruction fine-tuning. Claude3.5 (Anthropic, 2024) has enhanced performance, accuracy, and excels at high-level understanding and reasoning, providing reliable and consistent performance. Qwen2.5 (Yang et al., 2024) has strong instruction following ability, improved coding and mathematical ability, can generate long text, understand structured data, and generate structured output. Deepseek-R1 (Guo et al., 2025) is efficient and flexible, and performs well on reasoning, coding, and mathematics tasks with lower training cost. At the same time, it can be applied to various application scenarios that require multi-modal processing such as image description generation. Farui (Alibaba Cloud, 2024) is a large-model-based AI legal advisor trained on professional legal data, capable of legal knowledge understanding, reasoning, and generation.
We apply 0-shot settings to GPT-4o (Hurst et al., 2024), LLaMA3.1 (Grattafiori et al., 2024), Claude3.5 (Anthropic, 2024), Qwen2.5 (Yang et al., 2024), DeepSeek-R1 (Guo et al., 2025) and Farui (Alibaba Cloud, 2024). Details of LLMs are provided in Appendix A.
5.2 Experiment Settings
Here we describe the implementation of the claim generation method used in our experiments. Note that all LLMs are replaceable in this method. For all models (GPT-4o, Claude 3.5, DeepSeek-R1, Farui, LLaMA3.1, and Qwen2.5), we used the official APIs or open-source checkpoints available via OpenAI, Anthropic, Deepseek, Aliyun, or Hugging Face. See Appendix B for full API URLs.
We adopt a zero-shot setting for all models evaluated in this study. Each model is prompted to generate the plaintiff’s legal claims based solely on the factual description of a given case. We provide the full factual context as input to the model, without any additional fine-tuning, demonstrations, or retrieval-based augmentation. The prompt used across all models is as follows: “Please generate the plaintiff’s claims based on the following facts.”
5.3 Metrics
When constructing our dataset, we have considered including common NLG metrics such as ROUGE (Lin, 2004), BLEU (Papineni et al., 2002), and BERT SCORE (Zhang et al., 2020) to help users evaluate text quality. However, these metrics do not provide a good measure of the overall claim quality (Liu et al., 2016; Novikova et al., 2017; Chaganty et al., 2018; Sellam et al., 2020; Deutsch and Roth, 2021; Liu et al., 2023b). To better assess claim quality, we introduce two metrics: factuality and clarity for evaluating the claims. We employ LLM, specifically GPT-4o (Hurst et al., 2024), for scoring. GPT-4o evaluates claims using defined prompts for these metrics, as detailed in Appendix C. The robustness and reliability of our proposed metrics are detailed in Appendix D.2.
ROUGE (Lin, 2004) is a set of metrics used in the NLP task. We keep the results of ROUGE-1, ROUGE-2, and ROUGE-L. ROUGE-1 and ROUGE-2 refer to the overlap of unigram and bigram between the generated and reference documents, respectively. ROUGE-L is a Longest Common Subsequence (LCS) based statistics.
BLEU (Papineni et al., 2002) is used for automatic text-generation evaluation and correlates well with human judgment. We evaluate using BLEU-1, BLEU-2, and average scores from BLEU-1 through BLEU-4.
BERT SCORE (Zhang et al., 2020) measures similarity using contextual embeddings. We calculate the precision (p), recall (r), and f1-score to evaluate the information matching degree.
Factuality refers to the facts stated in the claims being truthful, accurate, and based on objectively existing circumstances.
Clarity means that claims should be specific, providing details such as the amount of compensation for losses and specifying the manner and scope of issuing an apology.
Total represents the average of factuality and clarity scores.
Method BLEU ROUGE BERT SCORE
B-1 B-2 B-N R-1 R-2 R-L p r f1
GPT-4o (Hurst et al., 2024) 19.07 7.25 7.85 48.98 22.40 43.78 76.06 65.86 70.46
LLaMA3.1 (Grattafiori et al., 2024) 16.95 5.64 6.82 43.14 19.12 36.05 69.38 70.58 69.64
Claude3.5 (Anthropic, 2024) 22.07 9.59 9.92 49.26 24.21 42.95 76.56 67.70 71.66
Qwen2.5 (Yang et al., 2024) 18.59 7.81 8.22 46.26 22.10 41.00 77.76 64.46 70.34
DeepSeek-R1 (Guo et al., 2025) 14.16 5.61 5.99 42.32 18.29 37.38 77.59 60.56 67.91
Farui (Alibaba Cloud, 2024) 20.81 8.54 9.08 48.60 23.10 42.55 76.28 66.45 70.85
Table 3: Automatic evaluation results on the ClaimGen-CN-test set using ROUGE, BLEU, and BERT SCORE. The best is bolded and the second best is underlined.
Method GPT-4o SCORE
Factuality Clarity Total
GPT-4o (Hurst et al., 2024) 48.31 56.23 52.27
LLaMA3.1 (Grattafiori et al., 2024) 51.32 55.18 53.25
Claude3.5 (Anthropic, 2024) 54.18 64.16 59.17
Qwen2.5 (Yang et al., 2024) 54.34 59.61 56.97
DeepSeek-R1 (Guo et al., 2025) 62.14 69.43 65.79
Farui (Alibaba Cloud, 2024) 42.85 46.28 44.56
Table 4: Evaluation results on the ClaimGen-CN-test set using GPT-4o. Total represents the average of factuality and clarity scores. The best is bolded and the second best is underlined.
5.4 Experimental Results
We report the performance of several state-of-the-art models on the ClaimGen-CN-test set using both automatic and human-aligned evaluations. The automatic metrics include BLEU, ROUGE, and BERT SCORE (Table 3), while GPT-4o-based human-aligned scores focus on factuality and clarity (Table 4).
Automatic Evaluation.
As shown in Table 3, Claude3.5 achieves the highest performance across most automatic metrics, including BLEU, ROUGE, and BERT SCORE f1. Farui ranks second in BLEU and BERT SCORE f1, while GPT-4o leads in ROUGE-L but is outperformed by Claude3.5 and Qwen2.5 on other metrics. These results suggest that Claude3.5 generates outputs with higher n-gram overlap and semantic similarity to the references, as captured by both lexical and contextual metrics.
GPT-4o Evaluation.
To better understand the alignment between generated claims and human judgments, we use GPT-4o to evaluate outputs based on Factuality and Clarity (Table 4). DeepSeek-R1 significantly outperforms other models with a total score of 65.79, showing the best factual accuracy and clarity. Claude3.5 again demonstrates strong performance with the second-best total score, while Qwen2.5 and LLaMA3.1 achieve moderate scores. While GPT-4o performs well on automatic metrics, its scores in GPT-4o-based evaluations are lower than those of several other models.
Overall, Claude3.5 stands out in terms of automatic metrics, whereas DeepSeek-R1 excels in GPT-4o evaluation. This discrepancy underscores the importance of incorporating human-centric evaluation to complement automatic metrics, especially for complex legal text generation tasks.
Figure 2: The claim generation of a given case. The purple parts indicate clarity mistakes in prediction, while the pink parts are factual mistakes in prediction. Green checkmarks indicate acceptable outputs, while red crosses mark outputs with factual or clarity errors.
5.5 Human Evaluation
To assess the alignment between GPT-4o evaluations and human judgments, we calculate the mean absolute error (MAE) and consistency score between GPT-4o and human annotations over 100 randomly sampled cases generated by DeepSeek-R1. We adopt Factuality and Clarity as evaluation dimensions. Three annotators are asked to assign scores on a 1-to-5 scale, where 1 denotes the worst and 5 the best.
Dimension MAE $\downarrow$ Consistency $\uparrow$
Factuality 0.19 81.05
Clarity 0.20 73.68
Table 5: MAE and consistency between GPT-4o scores and human annotations on 100 DeepSeek-generated samples.
As shown in Table 5, GPT-4o achieves a MAE of 0.19 and a consistency of 81.05 on the factuality dimension, and a MAE of 0.20 with 73.68 consistency on clarity. These results suggest that GPT-4o evaluations are closely aligned with human judgments, as evidenced by the low MAE and high consistency scores across both dimensions. In particular, the model demonstrates stronger agreement with human ratings on factuality, achieving a consistency of 81.05, compared to 73.68 on clarity. This indicates that while GPT-4o can serve as a trustworthy proxy for human annotation overall, its performance is more stable and reliable in assessing factual correctness than in evaluating clarity, where subjectivity may play a greater role.
A detailed analysis of human evaluation is provided in Appendix D.1, Appendix D.3 and Appendix D.4.
5.6 Case Study
Figure 2 illustrates a representative case from the ClaimGen-CN-test set, showcasing the comparative outputs of different large language models (LLMs) when generating legal claims. The factual background involves a plaintiff injured in a dispute, seeking compensation for subsequent treatment costs not covered by the defendant.
The ground truth specifies that the plaintiff requests the defendant to pay a total of ¥16,000 covering medical expenses, lost wages, transportation costs, and litigation expenses. Among the models evaluated, GPT-4o, Qwen2.5, and DeepSeek-R1 generate legally accurate and contextually faithful claims, aligning well with the facts and compensation details. In particular, GPT-4o excels in both factuality and clarity, offering a detailed yet concise articulation of the plaintiff’s request.
In contrast, models like Claude3.5 and LLaMA3.1 either misidentify the defendants or fabricate salary agreements, resulting in factual inconsistencies. Farui’s output, although structurally rich, suffers from vague and overly general phrasing, leading to reduced clarity and a failure to directly match the required compensation scope.
This case highlights the importance of grounding legal claim generation not only in accurate fact extraction but also in the precise mapping of legal entitlements, where minor deviations can significantly impact the validity of generated content.
6 Discussion
6.1 Error Analysis
Current models exhibit multidimensional deficiencies in legal document generation, primarily manifested through four core aspects: inadequate legal knowledge comprehension, disconnects in legal-mathematical logic, polarized deviations in claim generation, and systemic instability in output precision.
The following case examples are provided in Appendix E to supplement the error analysis discussed in this section. We include full input facts, ground truth and model outputs for all referenced instances: CaseID 1, 2, 3, 11, 93, 98, 110, and 116, in the order of appearance.
Lack of Legal Knowledge. We identify two common types of legal knowledge gaps that cause models to generate incorrect or incomplete claims. First, models often fail to reconstruct legally relevant facts from the case description. In CaseID 1, for instance, the model mistakenly assumed that interest began accruing a month too early. This shows the lack of ability to track and reason over event timelines based on contractual terms and payment records—an essential skill for accurate legal fact recognition. Second, models lack domain-specific legal knowledge needed for handling different types of cases. For example, in loan disputes, the validity of a claim often depends on understanding the sequence of events—loan issuance, due date, default, and interest—along with the corresponding legal rules. Without knowledge of such patterns, the model may generate structurally invalid or legally irrelevant claims.
[…]
Polarized Claim Generation. Models exhibit two opposite types of errors when generating claims: some produce redundant requests that are not supported by the facts, while others omit essential claims required by the legal context. In CaseID 3 and CaseID 11, GPT-4o and Qwen2.5 inserted mental damage compensation and interest claims that the plaintiff never requested, likely reflecting patterns overrepresented in debt-related training examples. On the other hand, Qwen2.5 and LLaMA3.1 omitted essential claims such as the confirmation of contract validity in CaseID 11, which is a legal prerequisite for pursuing the associated property transfer. These errors reflect two distinct but co-existing problems. First, some models tend to hallucinate legally implausible claims by defaulting to common case patterns rather than the actual fact statements. Second, others fail to identify necessary legal elements that are implied by the factual background or logically required by the cause of action. Consequently, claims that are either excessive or incomplete are produced, which reduces the factual and legal accuracy of the output.
Systemic Instability. LLaMA3.1 shows repeated output problems in two ways: (1) In CaseID 93, it copied and pasted over 180 consumer protection law articles without selecting the relevant ones, making the response bloated and meaningless; (2) In CaseID 98, 110, and 116, LLaMA3.1 repeated similar legal claims more than five times, showing a lack of control over output length and content. In practical terms, this hurts the clarity and coherence of generated claims, making it difficult for users to identify distinct legal requests. More importantly, these repeated and redundant outputs obscure the logical structure of the plaintiff’s position, potentially causing critical claims to be overlooked amid excessive or noisy output. This undermines the functional purpose of legal claim generation as a concise, actionable summary of the plaintiff’s demands.
6.2 Future Work in Legal Claim Generation
Our error analysis reveals deeper challenges in legal claim generation, pointing to several scenario-driven directions for future research. First, large–small model collaboration can improve factual grounding, such as using lightweight modules to identify key events or legal rules before invoking a larger model for structured claim generation. Second, long-chain reasoning techniques may help track complex legal timelines—e.g., loan issuance, default, and interest accrual—enhancing logical completeness. Third, reinforcement learning with legal-specific feedback can be used to optimize the claim generation process by designing task-specific reward functions. For instance, models can be penalized for producing claims that lack factual support, and rewarded for generating claims that follow legally valid reasoning paths. Such techniques should be developed in close alignment with real-world legal tasks to ensure reliability and practical value.
7 Conclusion
In this paper, we make a pioneering step in Legal AI by focusing on generating claims for civil litigation. Our dataset, ClaimGen-CN, covers a wide range of legal cases, offering a solid foundation for future Legal AI research and application. To evaluate model performance, we propose two dimensions, namely factuality and clarity, which are tailored to the specific requirements of legal text generation. Our error analysis reveals that current LLMs struggle with generating factually accurate and concise claims, suggesting key directions for future improvements.
Limitations
In this section, we discuss the limitations of our work:
$\bullet$ We only interact with the LLMs one round per time. The LLMs are capable of multi-round interaction, which may help the LLMs to better understand the claim generation task.
$\bullet$ The reliance on LLMs for claim generation raises questions about the transparency and interpretability of the generated claims. The decision-making process of these models is often opaque, which could lead to challenges in understanding and justifying the basis on which claims are generated.
$\bullet$ The exploration of legal claim generation tasks beyond the Chinese legal context remains limited.
Ethics Statement
The corpus we use is released by the Chinese government and has been anonymized wherever necessary [^5] Provisions of the Supreme People’s Court on Publication of Judgment Documents by the People’s Courts on the Internet 2016. http://gongbao.court.gov.cn/Details/415f49dd8baaa04b479d57af9616ef.html. Our dataset thus does not involve any personally private information. Besides, the corpus is in public domain and licensed for use within the legal scope [^6] See Article 3.3 of User Protocol of China Judgments Online. https://g.alicdn.com/onlineCourt/static/0.6.97/akan-wenshu-protocol.html.
With the increasing use of Legal AI for justice-related tasks, there is growing concern about its ethical implications, including the risk of biases and errors that could have serious repercussions. To address these concerns, it’s important to clarify that our work is an exploration of a new data source and the development of an algorithm, which is not meant to be immediately and directly used in practical settings. Our goal is to assist non-professionals by providing recommendations, not to make final decisions.
Measures for releasing the dataset: Our dataset is sourced from publicly available court documents and will undergo further validation to ensure its security. When releasing the dataset, we will provide usage limitations, appropriate scenarios, and guidelines. To minimize the risk of misuse, we will also implement access restrictions, including requiring applicants to provide their real identities.
Measures for future use: The proposed framework will include warning statements, such as adding the phrase “This answer may be incorrect and is for reference only” in the output. We will also insert declarations in the prompt instructing the model not to provide answers when confidence is low or requirements are not met, and we will consider additional validation after generation. The proposed framework will also advise users to seek professional legal advice, for instance, adding “Before making a final decision, you should consult a qualified lawyer” in the output.
Appendix A Details of LLMs Used in This Paper
As shown in Table 6, we employ six LLMs: GPT-4o (Hurst et al., 2024), LLaMA3.1 (Grattafiori et al., 2024), Claude3.5 (Anthropic, 2024), Qwen2.5 (Yang et al., 2024), DeepSeek-R1 (Guo et al., 2025) and Farui (Alibaba Cloud, 2024). The version of GPT-4o is gpt-4o-2024-08-06, the snapshot of gpt-4o from August 6th 2024. As for LLaMA3.1, we use the llama-3.1-8b-instruct version.The version of Claude3.5 that we use is Claude-3-5-sonnet-20241022, and the model was last updated in April 2024. We use Qwen2.5-7b-instruct as the version of Qwen, which is an instruction tuning model that performs well on reasoning, coding and mathematics tasks.The version of DeepSeek-R1 that we utilize is deepseek-reasoner, and Farui’s version is farui-plus. GPT-4o has a context window of 128,000 tokens. LLaMA3.1 (specifically the llama-3.1-8b-instruct version) has a context window of 131,072 tokens, surpassing that of Claude3.5 (200,000 tokens) and Qwen2.5 (128,000 tokens). DeepSeek-R1 features a context window of 64,000 tokens, and Farui has the smallest context window among them with 12,000 tokens. Each model’s version and characteristics are tailored to meet different application needs and scenarios.
Appendix B API Access URLs
Appendix C GPT-4o Scoring Prompts
The detailed prompts for two metrics are as follows.
For factuality.
“Factuality in claims requires that statements of fact be truthful, precise, and based on objectively existing circumstances. Please rate from 1 to 5 based on factual accuracy. Deduct points for errors by the plaintiff or defendant, discrepancies between factual descriptions and real scenarios, and inconsistencies between claims and actual events, whether adding or omitting details.”
For clarity.
“Clarity refers to statements that should be clear, concise, and unambiguous, avoiding vagueness, ambiguity, and unnecessary redundancy. Please rate from 1 to 5 based on clarity. Deduct points for repetitive or redundant reasoning sections and claims. Excessive additional information will result in a deduction of points.”
Appendix D Details about Human Annotation
D.1 Full Annotation Guidelines Provided to Annotators
To ensure consistency and fairness in human evaluation, we provided all annotators with a standardized set of instructions. Below, we report the full text of the annotation guidelines given to participants, in accordance with ethical best practices. This includes the task description, scoring rubric, and any disclaimers or risk notices if applicable.
Annotation Task Description
Annotators were instructed to evaluate each model-generated legal claim based on two criteria: Factuality and Clarity. Definitions of each are given below.
  • Factuality (1–5): Factuality in claims requires that statements of fact be truthful, precise, and based on objectively existing circumstances. Please rate from 1 to 5 based on factual accuracy. Deduct points for errors by the plaintiff or defendant, discrepancies between factual descriptions and real scenarios, and inconsistencies between claims and actual events, whether adding or omitting details.
  • Clarity (1–5): Clarity refers to statements that should be clear, concise, and unambiguous, avoiding vagueness, ambiguity, and unnecessary redundancy. Please rate from 1 to 5 based on clarity. Deduct points for repetitive or redundant reasoning sections and claims. Excessive additional information will result in a deduction of points.
Scoring Scale
Annotators were instructed to assign integer scores from 1 (very poor) to 5 (excellent), with intermediate scores indicating partial satisfaction of the criterion.
Disclaimer
Annotators were informed that the task involved reading real legal case descriptions and that no personally identifiable information (PII) was included. They were told they could opt out at any time. No personal risks were identified for participants in this task.
Figure 3 shows the interface provided to annotators for scoring legal claims generated by the model.
Figure 3: Interface used for human annotation.
D.2 Evaluation Robustness and Metric Validity
Meta-evaluation of the proposed metrics on SummEval (Fabbri et al., 2021) includes the Spearman correlation coefficient and Pearson correlation coefficient. The Spearman correlation coefficient and Pearson correlation coefficient between human and GPT-4o scores are respectively 0.5197 and 0.5248, which means a moderate correlation (Mukaka, 2012).
D.3 Annotation Protocol and Inter-Annotator Agreement
The inter-annotator agreement (IAA) between annotators and the correlation between human and GPT-4o scores are detailed as follows. The Fleiss’ Kappa among 3 annotators is 0.6823, indicating substantial agreement according to Landis and Koch’s interpretation scale (Landis and Koch, 1977), which validates the human evaluation process. The Spearman correlation coefficient between human and GPT-4o scores is 0.5197, and the Pearson correlation coefficient is 0.5248, both of which indicate a moderate correlation (Mukaka, 2012).
D.4 Metric Definitions and Scoring Procedure
We provide the formal definitions of the evaluation metrics used in Section 5.5, including the normalization process and the calculation of mean absolute error (MAE) and consistency.
Normalization. Given raw human and model scores $J_{i},P_{i}\in[s_{\min},s_{\max}]$, we normalize them to the $[0,1]$ interval as follows:
$\tilde{J}_{i}=\frac{J_{i}-s_{\min}}{s_{\max}-s_{\min}},\quad\tilde{P}_{i}=\frac{P_{i}-s_{\min}}{s_{\max}-s_{\min}}$
Metric Computation. Let $n$ be the number of evaluated samples. The metrics are computed as:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|\tilde{J}_{i}-\tilde{P}_{i}\right|$
$\text{Consistency}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left(\left|\tilde{J}_{i}-\tilde{P}_{i}\right|\leq\frac{\delta}{s_{\max}-s_{\min}}\right)$
In our experiments, we use a 5-point rating scale where $s_{\min}=1$, $s_{\max}=5$, and the tolerance threshold $\delta=1$. Thus, consistency corresponds to the proportion of samples where the absolute difference between model and human scores does not exceed 1.
Appendix E Case Examples for Error Analysis
We provide the full input and model-generated claims for the representative cases referenced in Section 6.1. For each case, we list the input facts (prompt”), the ground truth legal claim written by human experts (ground truth”), and the responses generated by each model. The selected examples include CaseID 1, 2, 31, 93, 98, 110, and 116, covering diverse types of factual patterns and model errors.
To facilitate close comparison, we highlight key elements in color: Blue text indicates critical factual expressions within the prompt. Green text marks the key elements in the ground truth claim that correspond to the legal intent. Red text is used to highlight incorrect or hallucinated content within the model outputs. Purple text highlights correct legal reasoning or factual reproduction by the models.
These annotated examples serve as the basis for our multi-dimensional error analysis in the main text, enabling a fine-grained understanding of how different models succeed or fail in legal claim generation.
Figure 4: CaseId 1.
Figure 5: CaseId 2.
Figure 6: CaseId 3.
Figure 7: CaseId 11-part1.
Figure 8: CaseId 11-part2.
Figure 9: CaseId 93-part1.
Figure 10: CaseId 93-part2.
Figure 11: CaseId 98.
Figure 12: CaseId 110-part1.
Figure 13: CaseId 110-part2.
Figure 14: CaseId 116-part1.
Figure 15: CaseId 116-part2.

### v2

@@
 Our findings highlight the limitations of the current models in factual precision and expressive clarity, pointing to the need for more targeted development in this domain.
-To encourage further exploration of this important task, we will make the dataset publicly available.
+ClaimGen-CN dataset is available at: https://github.com/JosieZhou00/ClaimGen-CN.
 ClaimGen-CN: A Large-scale Chinese Dataset for Legal Claim Generation
@@
 Our detailed error analysis reveals notable limitations in current models, particularly in terms of factual accuracy and linguistic clarity.
-  • To motivate other scholars to investigate this novel and important problem, we make the ClaimGen-CN dataset publicly available [^1] https://github.com/JosieZhou00/ClaimGen-CN.
 2 Related Work
@@
 Table 2: Comparison of legal datasets. – means the value is not accessible.
+Causes of Action Number
+Private Lending Disputes 37,781
+Motor Vehicle Traffic Accident Liability Disputes 24,025
+Sales Contract Disputes 17,487
+Financial Loan Contract Disputes 13,231
+Contract Disputes (General) 8,906
+Catering Service Contract Disputes 589
+Trademark Infringement Disputes 66
+Child Support Disputes in Cohabitation Relationships 70
+Bank Card Disputes 80
+Child Support Disputes 39
+Table 3: Cause of action distribution.
+Court Number
+Shanghai Pudong New District People’s Court 3,452
+Shanghai Minhang District People’s Court 2,329
+Shanghai Baoshan District People’s Court 1,620
+Shanghai Jing’an District People’s Court 1,470
+Shanghai Qingpu District People’s Court 1,390
+Ningbo Maritime Court 1
+Yueyang Intermediate People’s Court, Hunan Province 1
+Nantong Intermediate People’s Court, Jiangsu Province 1
+Lixian People’s Court 1
+Ankang Intermediate People’s Court, Shaanxi Province 1
+Table 4: Court distribution.
+Year Number
+2010 7
+2012 1
+2013 3
+2014 7
+2015 21
+2016 12,720
+2017 47,749
+2018 47,885
+2019 38,966
+2020 26,160
+2021 28,656
+2022 5,573
+Table 5: Case start year distribution.
+Statistic Fact Length Request Length Verdict Length
+Mean 353.4 134.3 605.8
+Std Dev 226.1 84.3 463.8
+Min 7.0 14.0 103.0
+Max 7,350.0 5,724.0 36,593.0
+Table 6: Text length distribution.
 4.2 Dataset Characteristics
@@
 Additionally, it shifts the focus from supporting judges to helping the non-expert public whose rights have been violated.
+Comprehensive.
+Beyond the general comparison, we further provide detailed distributions of ClaimGen-CN to demonstrate its comprehensiveness.
+As shown in Table 3, private lending disputes constitute the largest share, followed by motor vehicle traffic accident liability disputes and sales contract disputes, while a number of relatively rare categories such as trademark infringement and child support disputes also appear.
+Table 4 illustrates the court distribution, where cases are concentrated in Shanghai courts such as Pudong New District and Minhang District People’s Courts, though a small number of judgments come from courts in other provinces.
+Regarding temporal coverage, Table 5 shows that the dataset spans from 2010 to 2022, with a sharp increase starting in 2016 and peaking around 2017–2018 before declining in recent years.
+Finally, Table 6 reports the distribution of fact, request, and verdict text lengths, reflecting significant heterogeneity across different sections of judgments.
 5 Experiments
@@
 This discrepancy underscores the importance of incorporating human-centric evaluation to complement automatic metrics, especially for complex legal text generation tasks.
-Figure 2: The claim generation of a given case.
-The purple parts indicate clarity mistakes in prediction, while the pink parts are factual mistakes in prediction.
-Green checkmarks indicate acceptable outputs, while red crosses mark outputs with factual or clarity errors.
 5.5 Human Evaluation
@@
 This case highlights the importance of grounding legal claim generation not only in accurate fact extraction but also in the precise mapping of legal entitlements, where minor deviations can significantly impact the validity of generated content.
+Figure 2: The claim generation of a given case.
+The purple parts indicate clarity mistakes in prediction, while the pink parts are factual mistakes in prediction.
+Green checkmarks indicate acceptable outputs, while red crosses mark outputs with factual or clarity errors.
 6 Discussion
@@
 Such techniques should be developed in close alignment with real-world legal tasks to ensure reliability and practical value.
+Although our dataset is constructed based on the Chinese legal system, the overall data construction workflow and the evaluation design are not specific to China.
+With appropriate adjustments, especially considering the role of precedents in common law countries, our approach can be applied to other legal systems as well.
 7 Conclusion
@@
 https://g.alicdn.com/onlineCourt/static/0.6.97/akan-wenshu-protocol.html.
+Specifically, sensitive personal identifiers (such as names, identification numbers, and contact details) were removed or masked by the data providers before public release.
+Our research strictly used these desensitized datasets and did not involve any additional processing that could lead to re-identification.
+Beyond relying on officially anonymized and desensitized documents, we emphasize that our research activities did not introduce or generate any new sensitive information.
 With the increasing use of Legal AI for justice-related tasks, there is growing concern about its ethical implications, including the risk of biases and errors that could have serious repercussions.
@@
 Our goal is to assist non-professionals by providing recommendations, not to make final decisions.
-Measures for releasing the dataset: Our dataset is sourced from publicly available court documents and will undergo further validation to ensure its security.
-When releasing the dataset, we will provide usage limitations, appropriate scenarios, and guidelines.
-To minimize the risk of misuse, we will also implement access restrictions, including requiring applicants to provide their real identities.
+Measures for releasing the dataset: The dataset was solely used for academic purposes such as model evaluation and error analysis.
+No attempts were made to re-identify individuals, and no system or output of this work has been, or will be, deployed in real-world legal practice.
+We emphasize that AI systems in legal contexts must be subject to strict human oversight, and our work is intended to support research and understanding rather than to replace human judgment.
+Any potential application of similar technologies should therefore incorporate comprehensive oversight mechanisms to prevent misuse.
 Measures for future use: The proposed framework will include warning statements, such as adding the phrase “This answer may be incorrect and is for reference only” in the output.
@@
 The proposed framework will also advise users to seek professional legal advice, for instance, adding “Before making a final decision, you should consult a qualified lawyer” in the output.
+Acknowledgements
+We would like to thank the anonymous reviewers and the Area Chair for their constructive feedback and insightful suggestions.
+We are also sincerely thankful to Chen Yanwen for suggesting the name of the dataset and for providing valuable comments and revisions.
 Appendix A Details of LLMs Used in This Paper